# Unleashing the Potential of Large Language Models as Prompt Optimizers: An Analogical Analysis with Gradient-based Model Optimizers

### v2

@@ 3 Our Prompt Optimizer: GPO
Gradient
form
Momentum
form
Prompt
variation
Prompt
@@
GPO P+M Relevance Cosine Generation
Table 4: Comparisons of GPO with existing LLM-based prompt optimizers, including APE Zhou et al. (2023), APO Pryzant et al. (2023), OPRO Yang et al. (2023), and PE2 Ye et al. (2023). “P” refers to prompt, “M” refers to performance, and “R” refers to reflection.
Comparison of LLM-Based Prompt Optimizers. Existing LLM-based prompt optimizers can be divided into two main classes according to the update direction. The first line of research, such as APO Pryzant et al. (2023) and PE2 Ye et al. (2023), leverages the reflection capability of LLMs to produce textual “gradients” as the update direction. Another line of research, such as OPRO Yang et al. (2023) and APE Zhou et al. (2023), directly uses task prompts to derive the update direction. Our approach is based on the systematic investigation of the update direction as well as the update method. In particular, we propose several novel designs for the meta-prompt: relevance-based trajectory as the update direction and decay-based constraint for edit distance in the update method. Table 4 presents a detailed comparison.
4 Experiments
4.1 Experimental Setup
Tasks and Datasets. We select datasets from three groups of tasks for the experiment: Big-Bench Hard (BBH) Suzgun et al. (2023) and GSM8K Cobbe et al. (2021) for complex reasoning tasks, MMLU Hendrycks et al. (2021) for knowledge-intensive tasks, and WSC Levesque et al. (2012) and WebNLG Gardent et al. (2017) for common NLP tasks. Due to computational limitations, we sample a subset from each dataset for the main experiment. In addition, we use the lite BBH benchmark in Section 2.3.1 for detailed analysis. Other details are presented in Appendix D.1.
Baselines. We select several representative methods for comparison, including existing LLM-based prompt optimizers and one adapted from gradient-based model optimizers. (1) SGDM Sutskever et al. (2013) is a momentum-based model optimizer. We adapt it for prompt optimization using the summarization-based trajectory and the editing-based refinement strategy. (2) APE Zhou et al. (2023) utilizes LLMs to generate semantically similar variants of task prompts and selects one with the best performance. (3) APO Pryzant et al. (2023) first uses reflection to obtain the gradient and then edits the task prompt accordingly. (4) OPRO Yang et al. (2023) incorporates historical prompts with their scores into the meta-prompt. (5) PE2 Ye et al. (2023) adds historical prompts and reflection into the meta-prompt and controls the edit distance with a fixed constraint. In addition, we consider the baseline without an instruction (“Empty”) and the instruction “Let’ think step by step.” from chain-of-thought prompting (“CoT”) Kojima et al. (2022) for performance comparison.
Evaluation Metrics. We report the average accuracy of all the subtasks for BBH and MMLU following Suzgun et al. (2023) and Hendrycks et al. (2021), accuracy for GSM8K following Cobbe et al. (2021), ROUGE-L Lin (2004) for WSC and WebNLG following Wang et al. (2022).
Implementation Details. For the task model, we use both the base model (i.e., Llama-2-7b) and the instruction-tuned models (i.e., Llama-2-7b-chat, GPT-3.5-turbo, and GPT-4). For the prompt optimizer, we use GPT-3.5-turbo and GPT-4. Unless otherwise specified, we use Llama-2-7b-chat as the task model and GPT-3.5-turbo as the prompt optimizer throughout experiments. We repeat all the experiments three times and report the average of the results. Other details are presented in Appendix D.2.
4.2 Main Results
@@
Empty 30.51 22.00 35.96 60.67 32.14
@@
29.91 24.00 36.36 59.33 31.11
SGDM 33.30 27.33 37.30 64.00 38.01
APE 32.94 25.00 37.53 62.00 36.49
APO 32.97 25.33 37.75 62.00 34.92
OPRO 33.29 26.67 37.88 63.33 37.89
PE2 33.43 25.33 38.15 62.67 39.10
GPO 35.43 28.33 39.14 65.33 42.51
Table 5: Performance comparison using only the task prompts obtained from different methods.
Task model Llama-2-7b
Llama-2-7b-chat
@@
Inst. + Demo.
Inst.
Inst. + Demo.
Empty 40.28 32.29 36.63
CoT 36.46 31.25 34.20
SGDM 42.19 40.63 35.77
APE 42.54 42.01 36.29
APO 42.19 40.34 36.29
OPRO 42.02 42.14 36.46
PE2 42.88 42.01 36.81
GPO 45.48 43.75 38.02
Table 6: Performance comparison under different evaluation settings. “Inst.” denotes instruction and “Demo.” denotes demonstration.
Table 5 and Table 6 present the results of different methods for prompt optimization across various tasks and evaluation settings.
@@
Second, under various evaluation settings for the lite BBH benchmark, it can be observed that GPO not only excels in the “Instruction” setting but also yields substantial gains in the “Instruction + Demonstration” setting for both the base model and the instruction-tuned variant. Even in the scenario that is challenging for baselines (i.e., Llama-2-7b-chat with “Instruction + Demonstration”), our approach still demonstrates strong improvement over the empty prompt. This showcases the versatility of our approach in both zero-shot and few-shot evaluation settings.
4.3 Detailed Analysis
@@
Prompt optimizer
Task model
Accuracy
(before / after)
GPT-3.5-turbo Llama-2-7b-chat 31.25 / 43.75
GPT-3.5-turbo 62.15 / 67.02
GPT-4 73.61 / 76.56
GPT-4 Llama-2-7b-chat 31.25 / 44.97
GPT-3.5-turbo 62.15 / 67.80
GPT-4 73.61 / 78.65
Table 7: The performance before/after prompt optimization with different models as the prompt optimizer and the task model.
To confirm the effectiveness of GPO across different models, we explore the impact of model selection. Specifically, for the prompt optimizer, we employ GPT-3.5-turbo and GPT-4, while for the task model, we utilize Llama-2-7b-chat, GPT-3.5-turbo and GPT-4. Table 7 presents the results on the lite BBH benchmark. In general, GPO demonstrates remarkable capabilities for prompt optimization across various scenarios, including strong-to-weak optimization, self-optimization, as well as weak-to-strong optimization. This indicates the versatility of our framework. In particular, GPT-4 can consistently find better task prompts than GPT-3.5-turbo, which suggests the need for a capable model as the prompt optimizer.
@@
LLM-based prompt optimization requires multiple interactions with the LLM. In this part, we investigate the efficiency of GPO from the optimization curve of the first 12 steps and token consumption. First, Figure 2(a) shows that on the movie recommendation dataset, GPO exhibits rapid enhancement of the task prompt in the early stage, followed by consistent performance improvement in the later stage of optimization. Second, as depicted in Figure 2(b), the average token consumption of GPO on the lite BBH benchmark is much lower than SGDM, APO, and PE2, and comparable to APE and OPRO. Since GPO only utilizes task prompts to derive the update direction and performs fine-grained control over the variation, it can achieve better performance with high efficiency.
5 Related Work
Our work is related to the following two directions.
Prompt Engineering and Optimization. Prompt engineering aims to find suitable prompts as the input for LLMs to perform various tasks. To reduce human efforts, researchers have explored automatic prompt optimization, which can be categorized into continuous and discrete optimization methods. Discrete methods Deng et al. (2022); Xu et al. (2022) directly optimize the natural language prompts through methods like reinforcement learning Deng et al. (2022); Zhang et al. (2023) and editing Xu et al. (2022); Prasad et al. (2023). In contrast, continuous methods perform optimization in the embedding space of LLMs, allowing for optimization through gradient Li and Liang (2021); Lester et al. (2021). We focus on discrete methods, especially LLM-based prompt optimizers.
LLM-based Prompt Optimizers. Due to the unprecedented capabilities of LLMs, recent work starts to utilize them as prompt optimizers. One line of work Guo et al. (2023); Yang and Li (2023) combines LLMs with evolutionary algorithms to perform prompt optimization. Another line of work Pryzant et al. (2023); Yang et al. (2023) aims to adapt concepts and techniques from gradient-based model optimizers (e.g., gradient Pryzant et al. (2023) and momentum Yang et al. (2023)) to LLM-based prompt optimizers. However, no comprehensive guidelines exist for using LLMs as prompt optimizers. We aim to tackle this with a systematic investigation, which is conducted by analogy with gradient-based model optimizers.
6 Conclusion
We present GPO, a novel gradient-inspired LLM-based prompt optimizer. It utilizes LLMs to automatically optimize prompts, drawing inspiration from gradient-based model optimization techniques. By identifying the two crucial aspects of update direction and update method in model optimization, we enhance prompt optimizers by incorporating concepts from gradient-based model optimizers. Through extensive experiments, GPO demonstrates remarkable capabilities for prompt optimization across diverse tasks, models, and evaluation settings. Moreover, it surpasses competitive baselines while consuming fewer tokens.
In future work, we will explore more guidelines to further enhance the effectiveness and efficiency of LLM-based prompt optimizers.
7 Limitations
In this work, we improve the design of LLM-based prompt optimizers by drawing an analogy with the most widely used gradient-based optimizers in model optimization (e.g., gradient descent). More advanced model optimizers like Newton’s method Boyd and Vandenberghe (2014) and the corresponding improvement for meta-prompts remain to be investigated. In addition, due to the computational and budget limitations, we only conduct experiments on representative tasks and models.
Appendix A Extended Analysis
A.1 The Convergence of LLM-based Prompt Optimization
Figure 3: Optimization curve on Navigate dataset.
Figure 4: Optimization curve on Movie Recommendation dataset.
Figure 5: Optimization curve on Object Counting dataset.
Figure 6: Optimization curve on Word Sorting dataset.
@@ A.1 The Convergence of LLM-based Prompt Optimization
The convergence of LLM-based prompt optimization occurs when performance stabilizes at a plateau during the iterative process. Here, we illustrate the entire optimization curve on the lite BBH benchmark in Figure 3, Figure 4, Figure 5, and Figure 6. It can be observed that all the prompt optimization methods tend to reach a stable state within 12 steps. Therefore, our experimental settings (i.e., 3 epochs) are sufficient to demonstrate the effectiveness of GPO.
A.2 The Impact of Initial Prompts

### v3

@@ Introduction
Currect
information
Trajectory
Edit
distance
@@
GPO P+M Relevance Decay Generation
Table 4: Comparisons of GPO with existing LLM-based prompt optimizers. “P” refers to prompt, “M” refers to performance, and “R” refers to reflection.
@@
STEM Human. Social. Other Average
Empty 30.51 22.00 31.05 36.54 41.75 37.20 35.96 60.67 32.14
@@
29.91 24.00 32.53 37.05 41.05 36.94 36.36 59.33 31.11
SGDM 33.30 27.33 32.88 38.36 41.88 38.02 37.20 64.00 38.01
APE 32.94 25.00 33.51 38.69 42.02 37.96 37.50 62.00 36.49
APO 32.97 25.33 33.17 37.94 44.94 38.23 37.71 62.00 34.92
OPRO 33.29 26.67 34.76 38.72 43.55 37.11 38.05 63.33 37.89
PE2 33.43 25.33 33.77 37.95 44.80 38.25 38.07 62.67 39.10
GPO 35.43 28.33 35.00 38.84 46.61 38.60 39.14 65.33 42.51
Table 5: Performance comparison using only the task prompts obtained from different methods. “Human.” and “Social.” denote the datasets classified as Humanities and Social Science in MMLU.
In this section, we present our novel gradient-inspired LLM-based prompt optimizer called GPO, which leverages the insights gained from our systematic study. GPO adopts an iterative prompt optimization framework. At each step, it first retrieves relevant prompts from the optimization trajectory as the update direction. Then, it utilizes the generation-based refinement strategy to perform the update, while controlling the edit distance through a cosine-based decay strategy.
Iterative Prompt Optimization.
GPO performs prompt optimization through a multi-step iterative process. At each step, it first generates multiple candidate task prompts based on a meta-prompt. The meta-prompt serves as the input that guides the LLM in optimizing its prompts. Subsequently, we select the task prompt with the best performance for the next iteration. This iterative process continues until either the performance of the task prompt reaches a plateau or the predefined maximum number of optimization steps is reached.
The Design of Meta-Prompt.
As the input to the LLM, our meta-prompt consists of two key components: update direction and update method.
$\bullet$ Update direction. To determine the update direction, we leverage the retrieval-based optimization trajectory. This trajectory consists of past task prompts, along with their model performance. They are selected using a relevance-based strategy and sorted in ascending order based on their similarity to the current prompt.
$\bullet$ Update method. After the update direction is determined, we further employ the generation-based refinement strategy to update the task prompt. Specifically, we present the trajectory in the format of demonstrations in the meta-prompt. Then, the LLM can follow these demonstrations to generate a new task prompt via in-context learning. Additionally, we implement the cosine-based decay strategy to control the edit distance between task prompts at consecutive iterations, ensuring gradual and controllable changes.
Furthermore, we enhance the meta-prompt by incorporating a few task examples. These examples provide additional context to aid the LLM in understanding the task effectively.
Comparison of LLM-Based Prompt Optimizers.
Existing LLM-based prompt optimizers can be divided into two main classes by their update direction. Some work (i.e., APO (Pryzant et al. 2023) and PE2 (Ye et al. 2023)) leverages the reflection capability of LLMs to produce textual “gradients” as the update direction, while others (i.e., OPRO (Yang et al. 2023) and APE (Zhou et al. 2023)) directly use task prompts as the update direction. Our approach is based on the systematic investigation of the update direction and the update method. In particular, we propose several novel designs for the meta-prompt: relevance-based trajectory as the update direction and decay-based constraint for edit distance in the update method. Table 4 presents a detailed comparison.
Experiments
Experimental Setup
Tasks and Datasets.
We select datasets from three groups of tasks for the experiment: Big-Bench Hard (BBH) (Suzgun et al. 2023) and GSM8K (Cobbe et al. 2021) for complex reasoning tasks, MMLU (Hendrycks et al. 2021) for knowledge-intensive tasks, and WSC (Levesque, Davis, and Morgenstern 2012) and WebNLG (Gardent et al. 2017) for common NLP tasks. Due to computational limitations, we sample a subset from each dataset for the main experiment. In addition, we use the lite BBH benchmark for detailed analysis.
Baselines.
We select several representative methods for comparison, including existing LLM-based prompt optimizers and one from gradient-based model optimizers. (1) SGDM (Sutskever et al. 2013) is a momentum-based model optimizer. We adapt it for prompt optimization using the summarization-based trajectory and the editing-based refinement strategy. (2) APE (Zhou et al. 2023) utilizes LLMs to generate semantically similar variants of task prompts and selects one with the best performance. (3) APO (Pryzant et al. 2023) first uses reflection to obtain the gradient and then edits the task prompt accordingly. (4) OPRO (Yang et al. 2023) incorporates historical prompts with their scores into the meta-prompt. (5) PE2 (Ye et al. 2023) adds historical prompts and reflection into the meta-prompt and controls the edit distance with a fixed constraint. In addition, we consider the baseline without an instruction (“Empty”) and the instruction “Let’ think step by step.” from chain-of-thought prompting (Kojima et al. 2022) for performance comparison.
Task model
Llama-
2-7b
Llama-
3-8b
Llama-2-
7b-chat
Llama-3-
8b-instruct
@@
I + D
I
I + D
I
I + D
Empty 40.28 60.42 32.29 36.63 48.26 54.86
CoT 36.46 51.39 31.25 34.20 52.43 54.34
SGDM 42.19 62.50 40.63 35.77 60.42 58.51
APE 42.54 61.28 42.01 36.29 59.43 58.51
APO 42.19 61.45 40.34 36.29 59.38 57.99
OPRO 42.02 62.68 42.14 36.46 62.33 59.02
PE2 42.88 63.20 42.01 36.81 62.67 58.68
GPO 45.48 65.11 43.75 38.02 63.89 61.11
Table 6: Performance comparison under different evaluation settings. “I” denotes instruction and “D” denotes demonstration.
Prompt
optimizer
GPT-3.5-turbo GPT-4
Task
model
Baichuan2-
7b-chat
Llama-2-
7b-chat
Llama-2-
13b-chat
GPT-3.5-
turbo
GPT-4
Baichuan2-
7b-chat
Llama-2-
7b-chat
Llama-2-
13b-chat
GPT-3.5-
turbo
GPT-4
Empty 15.75 32.29 40.97 60.48 72.87 15.75 32.29 40.97 60.48 72.87
CoT 19.45 31.25 40.28 62.15 73.61 19.45 31.25 40.28 62.15 73.61
SGDM 20.87 40.63 42.24 64.18 75.23 21.85 40.41 42.03 64.55 75.82
APE 20.61 42.01 41.65 63.63 74.71 20.29 39.98 42.96 64.38 74.13
APO 19.97 40.34 41.79 63.96 74.13 22.09 39.56 43.67 64.55 74.56
OPRO 19.86 42.14 42.89 64.56 75.06 21.28 41.77 43.35 64.91 76.67
PE2 21.11 42.01 42.68 64.95 75.27 22.43 42.03 44.91 65.23 76.55
GPO 23.35 43.75 44.83 67.02 76.56 25.34 44.97 46.17 67.80 78.65
Table 7: Performance of different prompt optimization methods with various models.
Evaluation Metrics.
We report the average accuracy of all the subtasks for BBH and MMLU following Suzgun et al. (2023) and Hendrycks et al. (2021), accuracy for GSM8K following Cobbe et al. (2021), ROUGE-L for WSC and WebNLG following Wang et al. (2022b).
Implementation Details.
We use both the base model (i.e., Llama-2-7b and Llama-3-8b) and the instruction-tuned models (i.e., Baichuan2-7b-chat, Llama-2-7b-chat, Llama-2-13b-chat, Llama-3-8b-instruct (Dubey et al. 2024), GPT-3.5-turbo, and GPT-4) as the task model. For the prompt optimizer, we use GPT-3.5-turbo and GPT-4. Unless otherwise specified, we use Llama-2-7b-chat as task model and GPT-3.5-turbo as prompt optimizer throughout experiments. We repeat all the experiments three times and report the average of the results.
Main Results
Table 5 and 6 show the results of different methods for prompt optimization across various tasks and evaluation settings.
@@
Second, under various evaluation settings for the lite BBH benchmark, it can be observed that GPO not only excels in the “Instruction” setting but also yields substantial gains in the “Instruction + Demonstration” setting for both the base model and the instruction-tuned variant. Even in the scenario that is challenging for baselines (i.e., Llama-2-7b-chat with “Instruction + Demonstration”), our approach still demonstrates strong improvement. This showcases the versatility of our approach in both zero-shot and few-shot evaluation settings.
Detailed Analysis
@@
To confirm the effectiveness of GPO across different models, we explore the impact of different model combinations compared with other baseline methods. Table 7 presents the results on the lite BBH benchmark. In general, GPO demonstrates remarkable capabilities for prompt optimization across various scenarios, including strong-to-weak optimization, self-optimization, and weak-to-strong optimization. This indicates the versatility of our framework. Notably, GPT-4 can consistently find better task prompts than GPT-3.5-turbo, which suggests the need for a capable model as the prompt optimizer.
@@
LLM-based prompt optimization requires multiple interactions with the LLM. In this part, we investigate the efficiency of LLM-based prompt optimizers by examining the optimization curve over the first 12 steps. Figure 3 shows that on the movie recommendation dataset, compared to other methods, GPO demonstrates rapid enhancement of the task prompt in the early stage, followed by steady and consistent performance improvement in the later stage of optimization. Since GPO only utilizes task prompts to derive the update direction and performs fine-grained control over the variation, it can achieve better performance with high efficiency.
Conclusion and Discussion
In this paper, we conduct a systematic analogy between gradient-based model optimizers and LLM-based prompt optimizers. Based on existing work and our newly proposed approach, we conduct an experimental analysis of the two key factors (i.e., update direction and update method) to determine the best configuration. According to this configuration, we further propose a novel prompt optimization framework GPO. At each step, it retrieves relevant prompts from the optimization trajectory as the update direction. Then, it utilizes the generation-based refinement strategy to perform the update, while controlling the edit distance through a cosine-based decay strategy. Extensive experiments demonstrate the effectiveness and efficiency of GPO.
One limitation of our work is that we only draw an analogy with the most widely used gradient-based optimizers. More advanced model optimizers like Newton’s method and its application to meta-prompts remain to be investigated. Additionally, our approach relies on textual update directions, future research could explore more direct and fine-grained numerical update signal methods (Nie et al. 2024).
Acknowledgements
This work was partially supported by National Natural Science Foundation of China under Grant No. 92470205 and 62222215. Xin Zhao is the corresponding author.
A Additional Details for the Experimental Setup
In this part, we detail the setup for the analogical analysis and experiment.
A.1 Additional Details for the Setup of Analogical Analysis
Tasks and Metrics. Following Yang et al. (2023), we utilize Big-Bench Hard (BBH) (Suzgun et al. 2023) for evaluation. BBH is a suite of 23 challenging BIG-Bench tasks (Srivastava et al. 2022) that covers various kinds of reasoning capabilities. Due to the constraints of computational resources, we select a dataset from each type of task in BBH to create a lite benchmark for our analysis: i) Navigate (Binary choice); ii) Movie Recommendation (Multiple choice); iii) Object Counting (Numeric response); iv) Word Sorting (Free response). For each dataset, we split it into train/valid/test sets with a ratio of 2:2:6. Following Suzgun et al. (2023), we adopt the exact match as the metric for performance evaluation.
Implementation Details. Following Crispino et al. (2023), we select Llama-2-7b-chat as the task model and set its temperature to 0 to make the output as deterministic as possible. For the prompt optimizer, we employ GPT-3.5-turbo, which is the underlying model of ChatGPT. We set its temperature to 1.0 to encourage the generation to be more diverse. To help the prompt optimizer understand the downstream task, following Yang et al. (2023), we randomly sample 3 examples from the dataset and fill them into the meta-prompt of the prompt optimizer. In the process of optimization, we take the instruction “Let’s think step by step.” as the initial prompt and insert it at the end of the question to obtain better performance following Suzgun et al. (2023). At each step, we first prompt the optimizer to generate 8 candidate task prompts, and then select the best-performing one as the task prompt for the next iteration following Yang et al. (2023); Ye et al. (2023). The optimization process lasts for at most 3 epochs with a batch size of 8. We repeat all the experiments three times and report the average of the results. The meta-prompts we used are detailed in Appendix F.1 Analogical Analysis.
A.2 Additional Details for the Setup of Experiment
The Statistics of Datasets
As mentioned in the main body, we sample a subset of the dataset for efficient evaluation. For BBH and MMLU, we split the entire dataset into training, validation, and test sets with a ratio of 2:2:6. For GSM8k and WebNLG, we randomly sample 100 examples as the training set, 100 for the validation set, and 300 for the test set. For WSC, we randomly sample 50 examples as the training set, 50 for the validation set, and 150 for the test set.
Implementation Details
For the task model, we use both the base model (i.e., Llama-2-7b) and the instruction-tuned models (i.e., Llama-2-7b-chat, GPT-3.5-turbo, and GPT-4). The temperature is 0. For the prompt optimizer, we use GPT-3.5-turbo and GPT-4. The temperature is 1.0. Unless otherwise specified, we use Llama-2-7b-chat as the task model and GPT-3.5-turbo as the prompt optimizer throughout experiments. In the meta-prompt, we include 3 task examples and 7 historical task prompts. For the initial task prompt, we use the original ones from Hendrycks et al. (2021) for MMLU, “Let’s think step by step.” from Kojima et al. (2022) for GSM8K and BBH, and “Let’s solve the problem.” from Yang et al. (2023) for WSC and WebNLG. At each step, the optimizer generates 8 candidates, and the best-performing one is selected. The optimization lasts for at most 3 epochs with a batch size of 8. We repeat all the experiments three times and report the average of the results. The meta-prompts are detailed in Appendix F.2 Experiment.
B Extended Analysis
Method Prompt Accuracy
MultiArith AQUA
Zero-shot CoT Let’s think step by step. 58.00 19.69
OPRO
Take a deep breath and work on this problem step-by-step.
59.33 20.47
GPO
Thoroughly strategize and meticulously plan every single
aspect of your approach, leaving no room for oversight,
and give utmost attention to even the smallest details,
with the goal of achieving exceptional results and excelling.
62.33 24.80
B.1 The Convergence of LLM-based Prompt Optimization
(a) Optimization curve on Navigate dataset.
(b) Optimization curve on Movie Recommendation dataset.
(c) Optimization curve on Object Counting dataset.
(d) Optimization curve on Word Sorting dataset.
@@ B.1 The Convergence of LLM-based Prompt Optimization
The convergence of LLM-based prompt optimization occurs when performance stabilizes at a plateau during the iterative process. Here, we illustrate the entire optimization curve on the lite BBH benchmark in Figure 4(a), Figure 4(b), Figure 4(c), and Figure 4(d). It can be observed that all the prompt optimization methods tend to reach a stable state within 12 steps. Therefore, our experimental settings (i.e., 3 epochs) are sufficient to demonstrate the effectiveness of GPO.
B.2 The Generalizability of Optimized Prompts
To assess the generalizability of optimized prompts, following Zhou et al. (2023), we evaluate the prompts optimized on GSM8k (Cobbe et al. 2021) by OPRO and GPO on two additional math reasoning benchmarks MultiArith (Roy and Roth 2016) and AQUA (Ling et al. 2017). Table 8 demonstrates that the optimized prompts by GPO can be generalized well on these two benchmarks, which perform better than those by OPRO.
B.3 The Impact of Initial Prompts